use substitution method to determine whether the system of linear equations has no solution, one solution,or many solutions
2x - y = 8
x + 4y = 13

Answer by mananth(16946)

2 x -1 y = 8 .............1
4 y = 13 -1 x ---------2
/ 4 =
y = 3.25 -0.25 x
Plug the value of y in (1)
2x-1(3.25-0.25x) =8
2x-3.25+ 0.25x=8
2x+0.25 x=8+3.25
2.25x=11.25
/2.25
x=5
Plug the value of x in (1)
2x-y=8
2*5-y=8
10-y=8
-y=-2
y=2
(5,2) is the solution